for an aquaes solution having cacl2(5M) and mgcl2(5m) specfic density is 1.14g/ml. find MOLARITY OF CL- + MALALITY OF CL-

Grade:11

2 Answers

Arun
3 years ago
1 mole of CaCl2 has 2 moles of Cl-
 
5 moles of CaCl2 has = 2x5 = 10 moles of Cl-
 
1 mole of MgCl2 has 2 moles of Cl-
 
5 moles of MgCl2 has = 2x5 = 10 moles of Cl-
 
Total number of moles of Cl- = 10 + 10
 = 20 moles
 
The concentration of Cl ions in the solution in terms of molarity
= 20M
Vikas TU
3 years ago
1molecular Cacl2 has 2 moles cl- 
= 5M CaCl2 = 2*5 = 10M cl- 
1 molecular of MgCl2 has 2 moles Cl- 
 = 5M MgCl2 = 2*5 = 10M cl- 
Total concentration of cl- = 20M